Use the following table to answer questions 7 through 12 The table shows the peices and the quantities consumed in Carnivore Country Suppose the base year is 2008 Also, suppose that 2008 is the year the typical consumption busket as determined, so the quantities consuomedin 2008 are the only quaretities needed to celculate the CP! in eck year Year Price of Beel Quseity of Beet Price of Pork Quantity of Park 2008 2009 2010 2.00 2.50 2.75 100 90 105 $1.00 0.90 1.00 120 130 7. What is the value of the basket in the base year? a.$300 b $333 C. $18.75 d. $459.25 e none of the above 8. What are the values of the CPI in 2008, 2009, and 2010, respectively? a. 100, 111, 139.6 b. 100, 109.2, 116 c. 100, 1133, 125 d. 835, 942, 100 e none of the above 9. What is the inflation rate for 20092 a. 0 percent b. 9.2 percent c. I1 percent d. 133 percent e. none of the above 10. What is the inflation rate for 2010 a. 0 percent b. 10.3 percent c. 11 percent d. 13.3 percent e none of the above 11. The table shows that the 2009 inflation rate is biased upward because of a bias due to the introduction of new goods b. bias due to unmeasured quality change c substitution bias d base-year bias. e. none of the above. 12. Suppose the base year is changed in the table from 2008 to 2010. Also, suppose that the typical consumption basket was now determined in 2010 (now use the 2010 consumption basket). What is the new value of the CFI in 2009? a 906 b. 1000 c. 114.7 d. 134.3 e. none of the above 13. Suppose your income rises from $19,000 to $31,000 while the CFI rises from 122 to 169. Your standard of iving has likely a. fallen b. risen. c stayed the same. d. You cant tell without knowing the base ycar
Answer #1

We have base year as 2008

Answer for Question 1)

Price of basket in base year =2*100+1*100=$300

Answer for Question 2)

CPI in 2008=Price of basket in 2008/Price of basket in 2008=100

CPI in 2009=Price of basket in 2009/Price of basket in 2008=(2.5*90+0.9*120)/300=333/300*100=111

CPI in 2010=Price of basket in 2010/Price of basket in 2008=(2.75*105+1*130)/300=139.59

Option A is correct

Answer for 3)

Inflation in 2009=(CPI 2009-CPI 2008)/CPI 2008=111-100/100=11%

Option C is correct

Answer for 4)

Inflation in 2010=(CPI 2010-CPI 2009)/CPI 2009=28.6/111=25.76%

Hence answer is none of the above